# BANG: Billion-Scale Approximate Nearest Neighbour Search using a Single GPU

### v2

@@ 5       $u^{*}_{i}$, $\mathcal{L}_{i}$leftarrow $s$
             $N_{i}$leftarrow FetchNeighbors($u^{*}_{i}$, $G$)
              // on CPU
            7 /* CPU sends nbrs $N_{i}$ to GPU */
@@ 15                   $\mathcal{D}_{i}[k]\leftarrow$ ParallelComputeDist($N^{\prime}_{i}[k],q_{i}$)
16
            $(\mathcal{D}^{\prime}_{i},\mathcal{N}^{\prime}_{i})\leftarrow$ ParallelMergeSort($\mathcal{D}_{i},N^{\prime}_{i}$)
             // on GPU
             $\mathcal{L}_{i}$leftarrow ParallelMerge($\mathcal{L}_{i},\mathcal{D}^{\prime}_{i},\mathcal{N}^{\prime}_{i}$)
              // on GPU: Merge $\mathcal{N}^{\prime}_{i}$ with worklist $\mathcal{L}_{i}$ to keep $t$ NNs at most
@@ 19      $\mathcal{K}_{i}\leftarrow\ k$ candidates nearest to $q_{i}$ in $\mathcal{L}_{i}$
      19 /* $k$-nearest neighbours available on GPU */
20
@@
BANG performs greedy search on a proximity graph [25, 21, 31]. Algorithm 2 describes BANG’s scheme for batched- query searches on a GPU. Since all queries in a batch are independent, each query search can run in a separate CUDA thread-block (Section II-A) independently (see Line 2), resulting in as many thread-blocks as queries. It utilizes the GPU’s massive parallelism to maximize throughput, as measured by Queries Per Second (QPS). BANG addresses the challenge of the entire graph index not fitting on GPU efficiently by keeping the index on CPU and selectively retrieving from the CPU the neighbours of the visited nodes (i.e. candidate nodes) during query search iteration (Lines 2 and 2). This strategy ensures effective handling of billion-scale datasets on a single GPU, enhancing the overall performance of large-scale ANNS.
@@
Node Cluster PQ
ID ID Distance
0 6 6273
1 9 3893
2 7 3233
3 8 3433
4 1 3973
5 11 2113
6 0 1033
7 2 1853
8 4 433
9 5 673
10 3 73
11 10 593
Iteration Candidate Exact
Distance
0 6 1033
1 8 433
2 7 1853
3 2 3233
4 5 2133
5 9 673
6 11 593
7 10 73
@@
Figure 1 shows the complete workflow of BANG and the interaction between CPU and GPU. CPU maintains the graph and looks up its adjacency list to retrieve the neighbours of a node sent by the GPU. Similarly, the CPU maintains the complete base dataset as it cannot be fitted on the GPU, and the GPU performs the search using compressed vectors derived from the actual data vectors. Each step shown under the CUDA Cores section in Figure 1 runs in an individual CUDA thread-block (Section II-A) independently for each query. Figure 2 shows the workflow on a toy example, having 12 points in the base dataset, and its prebuilt Vamana graph index is shown in Figure 2(a). We use this running example to explain our technique, BANG.
@@
BANG addresses the challenges discussed in Section III by efficiently utilizing CPU, GPU, and PCIe bus hardware through a phased-execution approach. This strategy optimally balances ANNS workload across CPU and GPU while mitigating CPU-GPU data transfer bottlenecks by implementing prefetching techniques for asynchronous data transfer from CPU to GPU during specific phases. Further, by systematically dividing the ANNS activity into distinct steps, each with its own characteristics of work distribution and span, we maximize GPU parallelism, in contrast to other approaches [49, 21] that treat the entire search activity as a single block. That is, we design separate GPU kernels for each of these tasks within the search activity: (1) filtering out previously processed neighbours of the current node $u^{*}_{i}$ at Line 2, (2) computing distances of neighbours to the query $q_{i}$ at Line 2, (3) sorting the neighbours by computed distances at Line 2, and (4) merging the closest neighbours into the worklist at Line 2. Figure 2(c) shows these steps for each iteration of the query, where the GPU filters and sorts the neighbours by their distance to the query, merges it with $L_{i}$, and selects the first candidate node (e.g., 8 in iteration 1) to be visited next. For each individual kernel, we optimize the thread-block size (per query) to ensure maximum GPU occupancy, as a single sub-optimally configured block can significantly reduce performance. Further, this phased-execution strategy allows us to reduce GPU idle time by eagerly predicting the next candidate node $u^{*}_{i}$ immediately after neighbour distance calculation rather than waiting for the previous iteration to finish before communicating $u^{*}_{i}$ to the CPU.
@@
In addition to greedy search, BANG employs two additional steps which result in three stages that happen one after the other: 1 the distance table construction, 2 greedy search, and 3 re-ranking, as shown in Figure 1. Since the dataset cannot fit on GPU, it is compressed using the popular Product Quantization [28] technique. This technique partitions the $d$-dimensional dataset vector space into $m$ vector subspaces and performs independent $k$-means clustering of the dataset points(nodes) in each subspace, resulting in a centroid vector and ID for each cluster. Each dataset point is encoded using the cluster IDs belonging to the point in the respective $m$ vector subspaces. Eventually, the input $d$-dimensional vector is represented by a shorter $m$-dimensional vector. Figure 2(b) shows the cluster ID associated with each node. Since $m$=1 in the toy example, each 2-dimensional node vector is encoded into a single-byte cluster ID. We calculate the distances between centroid vectors and a given query vector within each subspace in a preprocessing step 1 and store them in a data structure PQ Distance Table (see Figure 2(b)). During the greedy search stage, distances between dataset points (in compressed form) and query vectors are calculated by summing the precomputed distances of the cluster IDs across all subspaces (i.e. available in PQ Distance table). Finally, after the greedy search process converges, BANG performs a re-ranking step 3 to refine the inaccurate set of $k$ nearest neighbours that might have resulted due to the use of compressed vectors. The re-ranking step calculates the exact distances (Figure 2(d)) of candidate nodes to the query point using the original data vectors, ensuring precise identification of the final list of $k$ nearest neighbours. (Note: The exact distances in Figure 2(d) and the corresponding PQ distances in Figure 2(b) happen to be identical for the respective nodes as it is a toy example. However, practically, they would be different.) For the toy example, the sorted final list of candidate nodes, ordered by their distances to the query node, includes 10, 8, 11, 9, 6, 7, 5, 2, resulting in the selection of the top 2 nearest neighbours, namely, 10 and 8. BANG performs both reranking and distance table construction on GPU. Note that, for the re-ranking step, only full vectors of selected nodes are sent to GPU, which are fewer (compared to sending the entire dataset) and can thus collectively fit into GPU memory for all queries.
@@
In the following subsections, we describe in detail how we parallelize and optimize the mentioned kernels and each step in Algorithm 2.
@@ IV-A Construction of $PQDistTable$ in Parallel
We maintain $PQDistTable$ as a contiguous linear array of size $(\rho\cdot m\cdot 256)$, where $\rho$ is the size of the query batch and $m$ is the number of subspaces with each subspace having 256 centroids. The number of centroids is as used in prior works [27, 25] that use Product Quantization [28]. We empirically determine $m=74$ for our setup, guided by the available GPU global memory (Section VI shows the ablation study with varying $m$ values). For example, in Figure 2(b), after compression, each input vector (2 bytes) gets compressed to a size of 1 byte using $m=1$ for the product quantization. Each thread-block handles one query, resulting in $\rho$ concurrent thread-blocks. Furthermore, for a given query $q\in\mathbb{R}^{d}$, the distance of each of the subvectors of the query $q_{s}\in\mathbb{R}^{d/m}$ from each of the 256 centroids of a subspace can be computed independently of others in parallel by the threads in a thread-block. Note that, for query $q$, distances of subvectors across $m$ subspaces are computed sequentially within a thread, ensuring an adequate workload per thread and constrained by thread-block size.
@@
$O((m\cdot subspace\_size)\cdot 256\cdot\rho)$. Owing to the parallelization scheme, the span of the algorithm is $O(m\cdot subspace\_size)=O(d)$, where d is the dataset dimension, and $subspace\_size$ is the size of the subspace (i.e., $d/m$).
@@ IV-B Handling Data Transfer Overheads
As outlined in Algorithm 2, the CPU transfers the neighbours (Line 2) to the search routine executing on the GPU for every candidate transmitted by the GPU (Line 2). The data transfer between the device (GPU) and the host (CPU) is time-consuming in contrast to the GPU’s tremendous processing power (the PCIe 4.0 bus connects the GPU and has a maximum transfer bandwidth of only 32 GB/s). Therefore, BANG transmits only the bare minimum information required in order to minimize data transfer overhead. Specifically, from device to host, the transfer includes a list of final approximate nearest neighbours after the search in Algorithm 2 converges and a candidate node for each query in every iteration (Line 2), while from host to device, it comprises a list of neighbouring nodes of candidate nodes (Line 2) and base vectors/coordinates of the candidates (not shown in the algorithm).
@@
Further, BANG hides data transfer latency with kernel computations using advanced CUDA features. To perform data transfers and kernel operations concurrently, CUDA provides asynchronous memcpy APIs and the idea of streams [42]; BANG makes use of these. During the search, neighbours of a given candidate are retrieved using CPU threads for all $\rho$ queries (see Line 2 in Algorithm 2). We leverage the efficient structure of the graph data, allowing sequential memory access of the node’s base vector and its neighbourhood list as they are placed next to each other on the host memory. Hence, immediately after the transfer of the neighbour list to the GPU in each search iteration, we strategically make an asynchronous transfer of the base dataset vectors for future use that are only required during the final re-ranking step on GPU (after the search in Algorithm 2 converges). As a result, the kernel execution engine and the copy engine of the GPU are kept occupied to achieve higher throughput. Thus, in our implementation, we aim to make all the data transfers asynchronous with cudaMemcpyAsync() and placing cudaStreamSynchronize() at appropriate places to honor the data dependencies.
IV-C Handling Processed Vertices: Bloom Filter
Given the nature of the Graph $G$ to establish long-range edges [25], Algorithm 2 (Line 2) may encounter the same node multiple times during the search, potentially impacting throughput and/or recall. The low throughput issue arises from the possibility of duplicate work during the search, particularly when fetching neighbours (Line 2) and computing their distances to query (Line 2), both of which are computationally intensive. Avoiding this redundancy presents an opportunity for throughput improvement. The low recall issue arises when not filtering out processed neighbours, leading to adding these neighbours to the worklist $\mathcal{L}_{i}$ multiple times (duplicate entries in worklist $\mathcal{L}_{i}$), potentially causing premature termination of the search. Further, the low recall issue has a stronger impact, as evidenced in our experiments where the recall diminishes to as low as 1/10th when processed vertices are not filtered out. Considering these factors, we keep track of the nodes processed during the traversal to ensure that a node is processed only once.
The most common methods for keeping track of the processed nodes are: i) setting a bit per each node and ii) using a set or a priority queue or a hash table data structure. While the first approach is fast, the memory footprint depends on the size of the dataset and the number of queries in a batch. For a billion-size graph and a query batch size of 10,000, it would require $(10^{9}\times 10^{4})$ bits or $\frac{10^{13}}{10^{9}\times 8}$ GB = 125 GB memory. This alone would be too large to fit on the GPU. The second approach has a small memory overhead since it only stores the vertices that are visited. However, concurrent dynamic data structures such as set, priority queue, or hash table are inefficient to maintain on GPU due to the massive multithreading and the changing memory requirements. These inherent characteristics result in underutilising GPU computing bandwidth as evidenced by previous works [49, 21] and thus are not conducive to GPU parallelization.
To optimize parallelism within the constraints of limited GPU memory, we employ the well-established Bloom filter [7] for its efficiency in approximate set-membership tests and minimal memory footprint, fully leveraging GPU parallelism. We use one Bloom filter per query as shown at Line 2 in Algorithm 2. Our Bloom filter uses an array of $z$ bools, where $z$ is determined using an estimate of the number of processed nodes for a query, a small tolerable false-positive rate, and the number of hash functions used. We use two FNV-1a hash functions [20], which are lightweight non-cryptographic hash functions often used to implement Bloom filters.
@@ IV-D Parallel Neighbour Distance Computation
When computing distances in GPU-based ANNS, summations are commonly conducted using reduction APIs from NVIDIA’s CUB [37] library, such as those in [21, 49], typically at a thread-block-level or warp-level. However, in our approach, a thread-block of size $t_{b}$ is subdivided into $g$ groups, each with $g_{size}\coloneqq\frac{t_{b}}{g}$ threads. These groups collaboratively compute the distance of the query from a neighbour, with $g_{size}$ threads summing up the partial distances of $m$ subspaces (with each thread processing a segment of size $\frac{m}{g_{size}}$). Each thread sequentially computes the sum of values in a segment using thread-local registers, avoiding synchronization. Finally, to efficiently add segment-wise sums using $g_{size}$ threads, we explore two approaches: (i) atomics, employing atomicAdd() for the final result; (ii) Sub-warp-level reductions, utilizing CUB [37] WarpReduce. Empirical tuning with $m=74$, $t_{b}=512$, and $g_{size}=8$ yields optimal performance in our experimental setup, with the second approach, utilizing sub-warp level reductions, marginally outperforming the first. Overall, our segmented approach outperforms standard alternatives like CUB WarpReduce (1.2×1.2$\times$1.2 × slower) and BlockReduce (4×4$\times$4 × slower) for warp-level and thread-block-level reductions.
This kernel constitutes a sizeable chunk of the total run time on billion-size datasets, accounting for approximately 38% on average. The kernel’s efficiency is limited by GPU global memory access latency arising from uncoalesced accesses to compressed vectors of various neighbours, a consequence of the irregular structure of the graph. To mitigate this irregularity, we explored graph reordering using the well-known Reverse Cuthill Mckee (RCM) algorithm [15, 19]. However, this approach did not yield significant improvements in locality, and so we avoided graph reordering.
@@
$O(R\cdot m\cdot\rho)$, where $R$ is the degree bound of the graph. Owing to the parallelization scheme, the span of the algorithm is $O(log~{}m)$.

### v4

@@ 5       $u^{*}_{i}$, $\mathcal{L}_{i}$leftarrow $s$
8             $N_{i}$leftarrow FetchNeighbors($u^{*}_{i}$, $G$) // on CPU
9            8 /* CPU sends nbrs $N_{i}$ to GPU */
@@ 17                   $\mathcal{D}_{i}[k]\leftarrow$ ParallelComputeDist($N^{\prime}_{i}[k],q_{i}$)
18
19            $(\mathcal{D}^{\prime}_{i},\mathcal{N}^{\prime}_{i})\leftarrow$ ParallelMergeSort($\mathcal{D}_{i},N^{\prime}_{i}$) // on GPU
20             $\mathcal{L}_{i}$leftarrow ParallelMerge($\mathcal{L}_{i},\mathcal{D}^{\prime}_{i},\mathcal{N}^{\prime}_{i}$) // on GPU: Merge $\mathcal{N}^{\prime}_{i}$ with worklist $\mathcal{L}_{i}$ to keep $t$ NNs at most
@@ 24      $\mathcal{K}_{i}\leftarrow\ k$ candidates nearest to $q_{i}$ in $\mathcal{L}_{i}$
25      24 /* $k$-nearest neighbours available on GPU */
@@
BANG performs greedy search on a proximity graph [29, 24, 36]. Algorithm 2 describes BANG’s scheme for batched- query searches on a GPU. Since all queries in a batch are independent, each query search can run in a separate CUDA thread-block (Section II-A) independently (see Line 2), resulting in as many thread-blocks as queries. It utilises the GPU’s massive parallelism to maximise throughput, as measured by Queries Per Second (QPS). BANG addresses the challenge of the entire graph index not fitting on GPU efficiently by keeping the index on CPU and selectively retrieving from the CPU the neighbours of the visited nodes (i.e. candidate nodes) during query search iteration (Lines 2 and 2). This strategy ensures effective handling of billion-scale datasets on a single GPU, enhancing the overall performance of large-scale ANNS.
@@
[] Refer to caption [] Node Cluster PQ ID ID Distance 0 6 6273 1 9 3893 2 7 3233 3 8 3433 4 1 3973 5 11 2113 6 0 1033 7 2 1853 8 4 433 9 5 673 10 3 73 11 10 593 [] Refer to caption [] Iteration Candidate Exact Distance 0 6 1033 1 8 433 2 7 1853 3 2 3233 4 5 2133 5 9 673 6 11 593 7 10 73
@@
Figure 1 shows the complete workflow of BANG and the interaction between CPU and GPU. CPU maintains the graph and looks up its adjacency list to retrieve the neighbours of a node sent by the GPU. Similarly, the CPU maintains the complete base dataset as it cannot be fitted on the GPU, and the GPU performs the search using compressed vectors derived from the actual data vectors. Each step shown under the CUDA Cores section in Figure 1 runs in an individual CUDA thread-block (Section II-A) independently for each query. Figure 2 shows the workflow on a toy example, having 12 points in the base dataset, and its prebuilt Vamana graph index is shown in Figure 2. We use this running example to explain our technique, BANG.
@@
BANG addresses the challenges discussed in Section III by efficiently utilizing CPU, GPU, and PCIe bus hardware through a phased-execution approach. This strategy optimally balances ANNS workload across CPU and GPU while mitigating CPU-GPU data transfer bottlenecks by implementing prefetching techniques for asynchronous data transfer from CPU to GPU during specific phases. Further, by systematically dividing the ANNS activity into distinct steps, each with its own characteristics of work distribution and span, we maximise GPU parallelism, in contrast to other approaches [67, 24] that treat the entire search activity as a single block (i.e. single kernel). That is, we design separate GPU kernels (as different phases) for each of these tasks within the search activity: (1) filtering out previously processed neighbours of the current node $u^{*}_{i}$ at Line 2, (2) computing distances of neighbours to the query $q_{i}$ at Line 2, (3) sorting the neighbours by computed distances at Line 2, and (4) merging the closest neighbours into the worklist at Line 2. Figure 2 shows these steps for each iteration of the query, where the GPU filters and sorts the neighbours by their distance to the query, merges it with $L_{i}$, and selects the first candidate node (e.g., 8 in iteration 1) to be visited next. For each individual kernel, we optimise the thread-block size (per query) to maximise GPU occupancy, as a single sub-optimally configured block can significantly reduce performance. Furthermore, this phased-execution strategy allows us to reduce GPU idle time by eagerly predicting the next candidate node $u^{*}_{i}$ immediately after neighbour distance calculation rather than waiting for the previous iteration to finish before communicating $u^{*}_{i}$ to the CPU.
@@
In addition to greedy search, BANG employs two additional steps which result in three stages that happen one after the other: 1 the distance table construction, 2 greedy search, and 3 re-ranking, as shown in Figure 1. Since the dataset cannot fit in GPU memory, it is compressed using the popular Product Quantization [32] technique. This technique partitions the $d$-dimensional dataset vector space into $m$ vector subspaces and performs independent $k$-means clustering of the dataset points(nodes) within each subspace, resulting in a centroid vector and ID for each cluster. Each dataset point is encoded using the cluster IDs belonging to the point in the respective $m$ vector subspaces. Eventually, the input $d$-dimensional vector is represented by a shorter $m$-dimensional vector. Figure 2 shows the cluster ID associated with each node. Since $m=1$ in the toy example, each 2-dimensional node vector is encoded into a single-byte cluster ID. We calculate the distances between centroid vectors and a given query vector within each subspace in a preprocessing step ( 1) and store them in a data structure PQ Distance Table (see Figure 2). During the greedy search stage ( 2), distances between dataset points (in compressed form) and query vectors are calculated by summing the precomputed distances of the cluster IDs across all subspaces (i.e. available in PQ Distance table). Finally, after the greedy search process converges, BANG performs a re-ranking step ( 3) to refine the inaccurate set of $k$ nearest neighbours that might have resulted due to the use of compressed vectors. The re-ranking step calculates the exact distances (Figure 2) of candidate nodes to the query point using the original data vectors, ensuring precise identification of the final list of $k$ nearest neighbours. (Note: The exact distances in Figure 2 and the corresponding PQ distances in Figure 2 happen to be identical for the respective nodes as it is a toy example. However, practically, they would be different.) For the toy example, the sorted final list of candidate nodes, ordered by their distances to the query node, includes 10, 8, 11, 9, 6, 7, 5, 2, resulting in the selection of the top 2 nearest neighbours, namely, 10 and 8. BANG performs both reranking and distance table construction on GPU. Note that, for the re-ranking step, only full vectors of selected nodes are sent to GPU, which are fewer (compared to sending the entire dataset) and can thus collectively fit into GPU memory for all queries.
@@
In the following subsections, we describe in detail how we parallelise and optimise the mentioned kernels and each step in Algorithm 2.
@@ IV-A Construction of $PQDistTable$ in Parallel
We maintain $PQDistTable$ as a contiguous linear array of size $(\rho\cdot m\cdot 256)$, where $\rho$ is the size of the query batch and $m$ is the number of subspaces with each subspace having 256 centroids. The number of centroids is as used in prior works [31, 29] that use Product Quantization [32]. We empirically determine $m=74$ for our setup, guided by the available GPU global memory (Section VI shows the ablation study with varying $m$ values). For example, in Figure 2, after compression, each input vector (2 bytes) gets compressed to a size of 1 byte using $m=1$ for the product quantization. Each thread-block handles one query, resulting in $\rho$ concurrent thread-blocks. Furthermore, for a given query $q\in\mathbb{R}^{d}$, the distance of each of the subvectors of the query $q_{s}\in\mathbb{R}^{d/m}$ from each of the 256 centroids of a subspace can be computed independently of others in parallel by the threads in a thread-block. Note that, for query $q$, distances of subvectors across $m$ subspaces are computed sequentially within a thread, ensuring an adequate workload per thread and constrained by thread-block size.
@@
$O((m\cdot subspace\_size)\cdot 256\cdot\rho)$. Owing to the parallelisation scheme, the span of the algorithm is $O(m\cdot subspace\_size)=O(d)$, where d is the dataset dimension, and $subspace\_size$ is the size of the subspace (i.e., $d/m$).
@@ IV-B Handling Data Transfer Overheads
As outlined in Algorithm 2, the CPU transfers the neighbours (Line 2) to the search routine executing on the GPU for every candidate transmitted by the GPU (Line 2). The data transfer between the device (GPU) and the host (CPU) is time-consuming in contrast to the GPU’s tremendous processing power (the PCIe 4.0 bus connects the GPU and has a maximum transfer bandwidth of only 32 GB/s). Therefore, BANG transmits only the bare minimum information required in order to minimise data transfer overhead. Specifically, from device to host, the transfer includes a list of final approximate nearest neighbours after the search in Algorithm 2 converges and a candidate node for each query in every iteration (Line 2), while from host to device, it comprises a list of neighbouring nodes of candidate nodes (Line 2) and base vectors/coordinates of the candidates (not shown in the algorithm).
@@
Further, BANG hides data transfer latency with kernel computations using advanced CUDA features. To perform data transfers and kernel operations concurrently, CUDA provides asynchronous memcpy APIs and the idea of streams [55]; BANG makes use of these. During the search, neighbours of a given candidate are retrieved using CPU threads for all $\rho$ queries (see Line 2 in Algorithm 2). We leverage the efficient structure of the graph data, allowing sequential memory access of the node’s base vector and its neighbourhood list as they are placed next to each other on the host memory. Hence, immediately after the transfer of the neighbour list to the GPU in each search iteration, we strategically make an asynchronous transfer of the base dataset vectors for future use that are only required during the final re-ranking step on GPU (after the search in Algorithm 2 converges). As a result, the kernel execution engine and the copy engine of the GPU are kept occupied to achieve higher throughput. Thus, in our implementation, we aim to make all the data transfers asynchronous with cudaMemcpyAsync() and placing cudaStreamSynchronize() at appropriate places to honour the data dependencies.
IV-C Parallel Filtering of Visited Neighbours
Algorithm 2 (Line 2) may encounter the same node multiple times during intermediate iterations of the search. Not filtering visited nodes may degrade throughput due to redundant distance computations (Line 2) and reduce recall by allowing repeated entries in the worklist $\mathcal{L}_{i}$, potentially displacing more promising nodes and causing premature termination of the search. Our experiments on the SIFT-1B dataset show that this leads to up to a 10× drop in recall compared to when visited nodes are filtered. In each search iteration, Line 2 returns up to $R\times\rho$ neighbours to the GPU, where $R$ is the graph’s degree and $\rho$ is the query batch size, necessitating efficient visited checks on this large set to avoid overhead. Tracking visited nodes on the GPU is challenging, as bitmaps incur high memory overhead for large datasets, and dynamic data structures such as priority queues or hash tables are inefficient on GPU as they underutilise its massive parallelism [67, 24]. Therefore, we adopt the well-known Bloom filter for its low memory footprint and suitability for GPU execution.
We use a separate Bloom filter per query and assign one query per thread block to enable fine-grained parallelism (Line 2). Parallel accesses to the Bloom filter do not require global synchronisation, as each query uses a separate filter within a thread block. In case of a race within a thread block, i.e., if thread $x$ sets an entry for a node and thread $y$ tests a different node that hashes to the same entry before $x$’s write completes, $y$ may see the entry as unset. However, this is safe, as it prevents incorrect inclusion of the unvisited node of $y$. Conversely, in the opposite case, $y$ may see the entry set by $x$ and wrongly mark its own node as visited. While this can be avoided by synchronizing reads before writes using a barrier like __syncthreads(), we skip synchronisation due to its overheads, and such cases are rare.
@@ IV-D Parallel Neighbour Distance Computation
When computing distances in GPU-based ANNS, summations are commonly conducted using reduction APIs from NVIDIA’s CUB [50] library, such as those in [24, 67], typically at a thread-block-level or warp-level. However, in our approach, a thread-block of size $t_{b}$ is subdivided into $g$ groups, each with $g_{size}\coloneqq\frac{t_{b}}{g}$ threads. These groups collaboratively compute the distance of the query from a neighbour, with $g_{size}$ threads summing up the partial distances of $m$ subspaces (with each thread processing a segment of size $\frac{m}{g_{size}}$). Each thread sequentially computes the sum of values in a segment using thread-local registers, avoiding synchronisation. Finally, to efficiently add segment-wise sums using $g_{size}$ threads, we explore two approaches: (i) atomics, employing atomicAdd() for the final result; (ii) Sub-warp-level reductions, utilising CUB [50] WarpReduce. Empirical tuning with $m=74$, $t_{b}=512$, and $g_{size}=8$ yields optimal performance in our experimental setup, with the second approach, utilizing sub-warp level reductions, marginally outperforming the first. Overall, our segmented approach outperforms standard alternatives like CUB WarpReduce (1.2×1.2$\times$1.2 × slower) and BlockReduce (4×4$\times$4 × slower) for warp-level and thread-block-level reductions.
This kernel constitutes a sizeable chunk of the total run time on billion-size datasets, accounting for approximately 38% on average. The kernel’s efficiency is limited by GPU global memory access latency arising from uncoalesced accesses to compressed vectors of various neighbours, a consequence of the irregular structure of the graph. To mitigate this irregularity, we explored graph reordering using the well-known Reverse Cuthill Mckee (RCM) algorithm [16, 22]. However, this approach did not yield significant improvements in locality, and so we do not perform graph reordering.
@@
$O(R\cdot m\cdot\rho)$, where $R$ is the degree bound of the graph. Owing to the parallelisation scheme, the span of the algorithm is $O(log~{}m)$.